Circular No. 5758

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 J. Blakeslee and J. Tonry, Massachusetts Institute of Technology, com-

municate: “We have measured the magnitudes of the progenitor of SN 1993J in the V and Kron-Cousins I bands. The images were taken on the night of 1992 May 1 with the 2.4-m telescope (+ 2048×2048 Loral CCD; scale 0′′.343/pixel) at MDM Observatory on Kitt Peak. The exposure times were 700 s in I and 300 s in V; the seeing was 1′′.05 in the I image and 1′′.14 in the V image. The night was not photometric, but we were able to calibrate the photometry to better than 3 percent by comparing the images to earlier ones, which did not encompass the supernova progenitor, taken under photometric conditions on the Kitt Peak 4-m telescope. Using an aperture of diameter 3′′.4 to minimize background contamination, we find: V = 20.88 ± 0.13, I = 19.51 ± 0.04. Correcting by the psf to an aperture of diameter 10′′, we find V = 20.79 ± 0.13, I = 19.44 ± 0.04; these are uncorrected for galactic extinction. The progenitor appears extended north-south as compared to nearby point sources. At least ten stars are visible within 10′′, with I-band magnitudes ranging from 1 to 3 mag fainter than the progenitor. Most of these are quite red. The nearest is an object with I = 20.4 situated 3′′.8 to the northwest.”

Further photoelectric photometry (cf. IAUC 5750) by H. Smith, E. Lansing, MI (S; forwarded by J. A. Mattei, AAVSO); R. Mutel and P. Sauerbrei, University of Iowa (MS; comparison stars from Corwin, IAUC 5742); and P. Prugniel, Observatoire de Haute Provence (P): Mar. 31.05 UT, V = 10.8 (S); Apr. 3.2, B = 11.98 ± 0.06, V = 11.69 ± 0.07, R = 11.23 ± 0.03, I = 10.96 ± 0.09 (MS); 4.2, B = 12.19 ± 0.04, V = 11.86 ± 0.06, R = 11.37 ± 0.05, I = 11.13 ± 0.06 (MS); 5.1, V = 11.86 ± 0.10, R = 11.34 ± 0.02, I = 11.03 ± 0.07 (MS); 6.5, B = 12.21 ± 0.07, V = 11.85 ± 0.04, R = 11.28 ± 0.08, I = 11.11 ± 0.04 (MS); 7.00, B_{c} = 12.36, B_{c}–R_{c} = +0.78 (P); 7.052, V = 11.8 (S); 8.00, B_{c} = 12.25, B_{c}–R_{c} = +0.79 (P); 9.00, B_{c} = 12.15, B_{c}–R_{c} = +0.79 (P). Mutel and Sauerbrei note that there is no evidence for variability in the reference stars.

SU TAURI Further visual magnitude estimates (cf. IAUC 5732): Mar. 31.14 UT,

11.3 (J. Bortle, Stormville, NY); Apr. 1.15, 11.8 (D. York, Abiquiu, NM); 4.11, 12.6 (York); 5.03, 13.3 (Bortle); 8.08, [13.5 (Bortle).

1993 April 10 Daniel W. E. Green